OpenStudy (anonymous):

Terry is going to purchase one item each from two quarter machines. The first quarter machine contains bracelets that are 4 different colors, including her favorite color, red. There are equal amounts of each color. The second quarter machine contains 1 sticky hand, 1 pencil topper, 2 erasers, 3 plastic rings, and 4 bouncing balls. What is the probability that Terry will get a red bracelet and a plastic ring by putting a quarter in each machine A. 1/11 B. 3/44 C. 1/22 D. 1/44

OpenStudy (paki):

1/5 * 4/11 = 4/55... your given options are wrong...

OpenStudy (anonymous):

@paki Actually, I mean this in a non-rude way, but your math is wrong. The odds are 3/11 and 1/4. Therefore, the correct answer is B. Because 3/11 * 1/4= 3/44.

OpenStudy (paki):

@Freflex yeah you are rite... thanks for correction... :-)